a woman is twice as old as her daughter twenty years ago she was four times as old as her daughter how old are they now
Answer by stanbon(75887)

a woman is twice as old as her daughter twenty years ago she was four times as old as her daughter how old are they
-----
Equations:
m = 2d
m-20 = 4(d-20)
----
Substitute for "m" and solve for "d":
2d-20 = 4d-80
2d = 60
daughter = 30 years (daughter's age now)
-----------------------
Solve for "m":
m = 2d = 60 years (mother's age now)
----------------
Cheers,
Stan H.
====================